OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY OF THE PREDISSOCIATED, QUASILIN-EAR B^{1}A′ STATE OF CHF

\underline{C. TAO}, C. MUKARAKATE AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; T. W. SCHMIDT AND S. H. KABLE, School of Chemistry, University of Sydney, NSW 2006, Australia.

We have recently observed transitions to the predissociated, quasilinear B^{1}A′ state of CHF using a fluorescence dip detected optical-optical double resonance technique.^{a} By exciting selected rovibronic states in intermediate A^{1}A′′ levels belonging to the progressions 2_{0}^{n}, 1_{0}^{1}2_{0}^{n}, and 2_{0}^{n}3_{0}^{1}, a variety of B^{1}A′ state levels were observed, extending to an energy of 7000 cm^{−1} above the B^{1}A′ state origin. In this talk, we will focus on the spectroscopy of the B^{1}A′ state and the excited state barrier to linearity. Experimental data will be compared with ab initio predictions.

aC. Tao, S. A. Reid, T. W. Schmidt, and S. H. Kable, J. Chem. Phys. 125, 051105 (2007).